Find the distance of an isolated positive point charge of 10 nC, in order for it to produce

an electric potential of 120 V.




1
Expert's answer
2022-02-23T12:11:07-0500

Answer

the distance of an isolated positive point charge of 10 nC, in order for it to produce

an electric potential of 120 V.

"r=\frac{Kq}{V}\\=\frac{9*10^9*10*10^{-9}}{120}\\=0.75m\\=75cm"